Suppose you need to standardize a sodium thiosulfate solution for a titration experiment. To do so, you will react it with a solution of iodine. You add a 1.00 mL aliquot of 0.0200 M KIO3 solution to a flask, followed by 3 mL of distilled water, 0.2 g of solid KI, and 1 mL H2SO4. You then titrate the solution with sodium thiosulfate solution in order to determine the exact concentration of Na2S2O3. The end point of the titration is reached after 0.86 mL of Na2S2O3 is dispensed from a microburet. What is the concentration of the standard sodium thiosulfate solution?

____M
Answer #1

KIO3 + 5 KI + 3 H2SO4     3 K2SO4 + 3 H2O + 3I2

Moles of KIO3 = (Molarity x Volume) / 1000 = 0.02 x 1 /1000 = 2 x 10-5 moles

Moles of KI = Mass/Molar mass = 0.2/166 = 1.20 x 10-3 moles  

One mole of KIO3 react with 5 moles of KI

2 x 10-5 moles of KIO3 react with 1.0 x 10-4 moles of KI

Therefore, KIO3 is the limiting reagent

Moles of I2 produced = 2 x 10-5 x 3 = 6 x 10-5 moles

2 Na2S2O3 + I2          Na2S4O6 + 2 NaI

One mole of I2 react with 2 moles of Na2S2O3

Therefore, 6.0 x 10-5 moles of I2 react with (2 x 6.0 x 10-5) = 1.2 x 10-4 moles of Na2S2O3

Molarity of standard sodium thiosulfate solution = [(1.2 x 10-4) x1000]/0.86 = 0.139 M